OpenStudy (anonymous):

Can you help me understand :)?! Thanks <3 A line passes through (4, 5) and (8, 9). Which equation best represents the line?

OpenStudy (mathstudent55):

There is the "two-point form of the equation of a line" which is exactly what you need: \(y - y_1 = \dfrac{y_2 - y_1}{x_2 - x_1}(x - x_1)\)

OpenStudy (anonymous):

A line passes through (4, 5) and (8, 9). Which equation best represents the line? A. y = x + 1 B. y = 1 C. y = 2x + 1 D. y = 5x + 8. Sorry here is the answers to.

OpenStudy (mathstudent55):

Select one of the points to by \((x_1, y_1)\) and the other point to be \((x_2, y_2) \). Then just plug in the values into to the equation above.

OpenStudy (mathstudent55):

Since you are given choices, you can test both points in each choice. The answer is the equation that is true for both points.

OpenStudy (anonymous):

I might be wrong but from looking at it im thinking it might be the 2nd one. which is y=1